OpenStudy (anonymous):

Explain in complete sentences how to use the distributive property to simplify the expression below 5(x + 2) write the simplified expression

OpenStudy (unklerhaukus):

\[a(b\star c)\] the distributive property says \[a(b\star c)=ab\star ab\]

OpenStudy (unklerhaukus):

so \[5(x+2)=5\times x+5\times2=\]